# Guiding Generative Models for Protein Design: Prompting, Steering and Aligning

### v1

@@ 1.2 Inference-time control
With the aim to dynamically inject additional knowledge into the GMPD’s context, Retrieval-augmented generation (RAG) has been successfully applied in different cases[56, 57, 58]. By drawing upon semantically related examples from large databases, RAG enables more informed sampling and can enrich the design process with functional or structural priors not explicitly encoded in the pretrained model. A recent example is Protriever[57], which introduces a retrieval-augmented protein language model that jointly learns to retrieve homologous sequences and model their fitness, integrating evolutionary context at inference time without explicit structural supervision.
@@
A more surgical form of intervention acts directly within the hidden states of the network. Activation steering manipulates internal representations, often within the residual stream, by injecting vectors that correspond to interpretable latent directions. Sparse autoencoders (SAEs) have been used to identify such interpretable features in protein language models, revealing latent dimensions correlated with properties like enzymatic activity, hydrophobicity, or thermostability [62, 63]. For example, Parsan et al. used SAE-derived features to bias structure predictions in ESMFold toward more hydrophobic conformations through feature steering[64], while Boxò et al. leveraged activity-associated features to steer ZymCTRL toward more active $\alpha$-amylases[65]. This approach provides a surgical way of shaping the trajectory of the model’s activations without altering inputs or outputs explicitly.
Control can also be applied at the level of output probabilities. Bayesian guidance reweighs the probability distribution encoded by GMPDs using Bayesian principles, effectively combining the model’s prior with external evidence or predictive scores. Such strategies have been applied in protein design, where sequence likelihoods are updated according to functional predictors or activity [66].
@@
Finally, Sampling Controls manipulate the stochasticity of the GMPD’s final output by, for example, manipulating "inference parameters" like temperature, top-k and top-p sampling, balancing diversity and fidelity, which is particularly important when sampling from vast sequence landscapes [67]. More advanced sampling techniques, such as beam search and Monte Carlo Tree Search (MCTS), allow to consider multiple GMPDs inference trajectories, selecting the optimal ones [e.g., 9, 68].
@@ 1.3 Conclusion and future prospects
Protein design operates at the intersection of evolutionary complexity and computational abstraction. Natural proteins are shaped by diverse and often competing forces: biochemical constraints, ecological pressures, and evolutionary contingencies; making it difficult to define a single, global “fitness vector”. Instead, the true fitness landscape is fluid, heterogeneous, and context-dependent. Combined with well-known dataset biases [e.g., see 19, 20, 22], these properties impose severe limitations on how much we can leverage the natural distribution of proteins to design and engineer proteins à la carte, with full control over the design process.
Here, we discussed two broad categories of methods, fine-tuning and inference-time control, which are arising as powerful tools to guide GMPDs towards desired regions of the learned protein distributions (e.g., those encoding target design properties), giving us tighter control over the protein design process.
However, a recurring limitation of both families is poor out-of-distribution (OOD) generalization. Inference-time control presupposes that GMPDs already encode rich structure over the relevant regions of sequence space; steering cannot along produce representations that the base model never learned. Fine tuning methods may allow for more flexibility, enabling to reshape the GMPD’s learned distribution by changing its parameters. However, many curated datasets used for supervised fine-tuning and in-silico scoring methods used for RL are based on the natural distribution of proteins, reflecting the same biases and evolutionary constraints we would like to move away from [e.g., 69, 18, 60].
In this regard, including physics-based scoring methods, like RoseTTA fold [70] and FoldX [71], in the RL fine-tuning process may provide a promising avenue, enabling to move away from models that exclusively infer information from the natural distribution of proteins. [72].
Moreover, as GMPDs are pushed to explore protein regions distant from the natural distribution of proteins, experimental testing providing reliable validation is paramount. In this regard, lab-in-the-loop frameworks combining GMPDs with experimental validation are a very promising avenue. Advances in laboratory automation [73] and techniques[74, 75] are reducing experimental bottlenecks, enabling faster experimental testing and unprecedented scale of screening.
A complementary strategy is to improve the data distribution seen during pretraining and downstream optimization: diversifying sequence sampling across the tree of life [76]. Other approaches introduce inductive biases, leveraging evolutionary context provided by Multiple Sequence Alignment to provide more efficient training [77] or sampling[22].
The field of generative modeling for protein research is advancing at an unprecedented pace. A multitude of techniques for the guidance and control of GMPDs has emerged in the past two years, demonstrating promise for the targeted engineering of proteins across diverse applications. As the field progresses, a central challenge will be extending these models beyond the natural sequence distribution—ensuring reliable performance in out-of-distribution (OOD) regimes and paving the way toward truly generalizable, controllable, and creative protein design.

### v2

@@ 1.2 Parameter-fixed steering
Alternatively, architectures such as BoltzGen[3] implement a conditional generative diffusion model with continuous guidance, in which encoded design conditions, such as binding-site specifications or structural constraints, are propagated throughout the denoising process to steer generation toward conformations consistent with the imposed design criteria [71]. With the aim to dynamically inject additional knowledge into the GM’s context, Retrieval-augmented generation (RAG) has been successfully applied in differentcases[39, 40, 41]. By drawing upon semantically related examples from large databases, RAG enables more informed sampling and can enrich the design process with functional or structural priors not explicitly encoded in the pretrained model. A recent example is Protriever[40], which introduces a retrieval-augmented protein language model that jointly learns to retrieve homologous sequences and model their fitness, integrating evolutionary context at inference time without explicit structural supervision.
Intervention at the hidden states.
@@
A different approach acts directly within the hidden states of the network, targeting the second stage. Activation steering manipulates internal representations, often within the residual stream, by injecting vectors that correspond to interpretable latent directions. Sparse autoencoders (SAEs) have been used to identify such interpretable features in protein language models, revealing latent dimensions correlated with properties like enzymatic activity, hydrophobicity, or thermostability [42, 43]. For example, Parsan et al. used SAE-derived features to bias structure predictions in ESMFold toward more hydrophobic conformations through feature steering[44], while Boxò et al. leveraged activity-associated features to steer ZymCTRL toward more active $\alpha$-amylases[72].
Intervention at the output stage and sampling.
@@
Finally, control can also be applied at the level of output. Bayesian guidance reweighs the probability distribution encoded by GMs using Bayesian principles, effectively combining the model’s prior with external evidence or predictive scores. Such strategies have been applied in protein design, where sequence likelihoods are updated according to functional predictors or activity [46]. Similarly, sampling Controls manipulate the stochasticity of the GM’s final output by, for example, manipulating "inference parameters" like temperature, top-k and top-p sampling, balancing diversity and fidelity, which is particularly important when sampling from vast sequence landscapes [48]. More advanced sampling techniques, such as beam search and Monte Carlo Tree Search (MCTS), allow to consider multiple GMs inference trajectories, selecting the optimal ones [e.g., 9, 47].
@@ 1.3 Conclusion and future prospects
Natural proteins are shaped by diverse and often competing forces: biochemical constraints, ecological pressure, and evolutionary mechanisms; making it difficult to define a single, global “fitness vector” that can be captured by GMs. Combined with well-known dataset biases [e.g., see 22, 23, 25], these properties impose severe limitations on how much we can leverage the natural distribution of proteins to design and engineer proteins à la carte, with full control over the design process.
Over the last two years, we have witnessed the emergence of several powerful methods for guiding GMs toward desired regions of the learned protein distribution (i.e., those encoding target properties for engineering). Here, we describe these methods and classify them into two broad categories: Parameter-Updating Alignment and Parameter-Fixed Steering. These strategies have already been used to bias protein generative models toward brighter fluorescent proteins (e.g., CreiLOV variants)[56], increased stability[32], reduced MHC-I epitope load[59], and improved target-binding designs, such as low-nanomolar EGFR inhibitors[33].
Despite this progress, several challenges still limit the immediate applicability of these approaches to fully autonomous protein engineering. A recurring issue across both categories is poor out-of-distribution (OOD) generalization. Parameter-Fixed Steering relies on the assumption that the pretrained model already encodes meaningful structure in the relevant regions of sequence space; consequently, steering cannot readily extrapolate beyond representations the base model has not learned. Parameter-Updating strategies offer greater flexibility by reshaping the learned distribution through fine-tuning, potentially extending coverage beyond pretraining. However, this flexibility comes at the cost of increased dependence on downstream data quality and optimization signals.
Many curated datasets used for supervised fine-tuning, as well as in silico scoring functions employed in reinforcement learning, remain rooted in the natural protein distribution and therefore inherit the same evolutionary biases and constraints that we ultimately attempt to overcome [e.g., 73, 21, 36]. In this context, lab-in-the-loop frameworks that iteratively combine generative models with experimental validation represent a particularly promising direction, as they could progressively enrich the data distribution [74]. Complementary efforts aim to improve pretraining and optimization data by diversifying sequence sampling across the tree of life [74] or introducing inductive biases that leverage evolutionary context through Multiple Sequence Alignments, enabling more efficient training [75] and sampling [25].
A second major bottleneck concerns the availability of accurate and reliable scoring metrics. For many properties of interest—such as catalytic activity—robust predictive oracles remain difficult to obtain. Incorporating physics-based scoring methods, such as Rosetta [76] and FoldX [77], into reinforcement learning pipelines may help ground optimization in physically informed principles and reduce reliance on purely data-driven signals.
More fundamentally, it remains unclear whether alignment and steering methods genuinely uncover novel functional solutions or primarily recombine patterns already present in the training data. Addressing this question will require systematic and controlled empirical comparisons across diverse tasks and evaluation settings, allowing the field to rigorously assess the relative strengths and limitations of Parameter-Updating and Parameter-Fixed approaches[60, 78]. As in many areas of biology, different strategies may prove optimal in different regimes—for example, de novo generation versus sequence optimization. Ultimately, advancing beyond the natural sequence distribution while maintaining reliability will be the defining challenge of the next phase of protein generative modeling, shaping the path toward truly controllable and new-to-nature protein design.